# Multipartite Entanglement Distribution in Quantum Networks using Subgraph Complementations

### v4

@@ IV-B Time for subgraph complementation
It is possible that $T(G,\mathcal{SC}_{sys})$ can be minimized by some $\mathcal{SC}_{sys}$ that is not of minimum size $\mathbb{c}_{2}(G)$. However, the case $d=\mathbb{c}_{2}(G)$ provides an upper bound $T(G,\mathcal{SC}_{sys})=\mathbb{c}_{2}(G)\cdot t(\overline{k})$ on distribution time, which we conjecture is a tight bound for most graph states.
Given a particular $\mathcal{SC}_{sys}$, we can also optimize the ordering. Note that we do not always have to do a complete edge reset, there are some edges that we could potentially keep, minimizing both the number of gates and time steps needed.
@@
If $c_{2}(G)$ is constant with respect to some class of graphs, then the graph states corresponding to that class could be efficiently distributed in a linear number of time steps since $t(\overline{k})\leq n$, so $T(G,\mathcal{SC}_{sys})=O(n)$ (from IV.1). It turns out we can characterize the class of graphs with a constant bounded value of $c_{2}(G)$. Formally, $\{G|\mathbb{c}_{2}(G)\leq d\}$ can be characterized by a finite number of forbidden induced subgraphs (i.e. there is no subgraph in $G$ isomorphic to any graph in some finite set) for all $d\in\mathbb{N}$ [10]. Dense structured graphs often have a bounded $d$, such as cliques ($d=1$) and complete bipartite graphs $(d=3)$. Specifically, the $\mathcal{SC}_{sys}$ with minimum $T(G,\mathcal{SC}_{sys})$ for cliques just corresponds to one $SC$ on all vertices. Similarly, for complete bipartite graphs with vertex partitions $P_{1},P_{2}$ the corresponding optimal subgraph complementation system is given by $\mathcal{SC}_{sys}=\{P_{1}\cup P_{2},P_{1},P_{2}\}$.
@@
Calculating the optimal number of time steps taken to distribute a graph state and the corresponding $\mathcal{SC}_{sys}$ appears to be a hard problem in general, but we observe some common characteristics of entanglement distribution using subgraph complementations. There is often a trade-off between $t(\overline{k})$ and $\mathbb{c}_{2}(G)$, for some $\mathcal{SC}_{sys}$, such that their product $T(G,\mathcal{SC}_{sys})=O(n)$ for structured or common classes of graphs. Intuitively, if larger subgraphs are complemented at each step, then fewer steps are needed to build an arbitrary structured graph, and vice-versa. We also note that greedy vertex cover algorithms provide good constant factor approximations for $\mathcal{SC}_{sys}$ w.r.t. $\mathbb{c}_{2}(G)$ [10], and hence also $T(G,\mathcal{SC}_{sys})$. Table I lists the asymptotic costs for several relevant classes of graphs.
TABLE I: Costs for distributing various graph states
@@
Sparse Graphs: $O(n)$ $O(1)$ $O(n)$
Path Graph $n-1$ $O(1)$ $O(n)$
Cycle Graph $n-2$ $O(1)$ $O(n)$
@@
Dense Graphs: $O(n)$
Complete Graph ($n$-GHZ state) $1$ $O(n)$ $O(n)$
Repeater Graph [12] $\leq 3$ $O(n)$ $O(n)$
@@
IV-C Entanglement of the graph state
An interesting property of $\mathbb{c}_{2}(G)$ is its relation to the Schmidt-rank $r_{G}(X)$, a common entanglement measure of pure states. For a graph state $\left|G\right\rangle$, $r_{G}(X)$ is defined over some bipartition $(X,V(G)\setminus X)$ of the vertices of $G$. The Schmidt-rank quantifies the number of superpositions of a pure state into factorizable states [13] —- a higher Schmidt-rank of a state implies it has higher entanglement. Define the average Schmidt-rank $\mathbb{E}_{r}(G)$ to be the expectation of $r$ for a uniformly chosen random subset of $V$.
Theorem IV.3.
@@ IV-C Entanglement of the graph state
Consequently, graph states with a high degree of entanglement will have large values of the subgraph complementation number $\mathbb{c}_{2}(G)$, and therefore may be more expensive to distribute.
V Comparison to prior work
Our approach utilizes $(n-1)$ EPR pairs, $O(\overline{k}n)$ bits of classical communication (CC), and a maximum of $n-1$ qubits at the central node, all of which improve on prior work [2, 3, 8]. Our approach requires one time step to distribute an arbitrary graph state, if we follow early work to implement our cost model [2, 16], which make the assumption that local gates at multiple nodes and the same Clifford operations on a single central node commute, so all the operations can be performed in one time step. We instead assume a gate-based model of computation, which implies $T(G,\mathcal{SC}_{sys})$ number of time steps is required. Other work [8] utilizes an approach with a central network node that takes $O(n)$ time steps if local operations are free and $O(|E|)=O(n^{2})$ time steps if not (accounting for local gates) to build a graph state corresponding to a dense graph. In the presence of memory decoherence, local operations are not free. The fidelity of the state is directly dependent on the number of time steps required. Our approach may also utilize $O(n^{2})$ number of gates and time steps to distribute certain dense graph states but not for most of the common classes since $T(G,\mathcal{SC}_{sys})=O(n)$ even for several classes of dense graphs. This presents a quadratic improvement in the number of time steps, gates required, and potential fidelity of the state in the presence of noise.
VI Incorporating Parallelization
VI-A Log depth GHZ state generation
We first describe a log depth circuit to generate a GHZ state, that is the same up to local rotations as the construction proposed by Cruz et al. [17]. This will be useful as a subroutine in reducing the time to create a star graph in steps 1 and 3 of the subgraph complementation protocol from III-A.
Let $n$ qubits be denoted by $q_{0},q_{1},q_{2}...q_{n-1}$. Begin with a GHZ state (star graph state with center $q_{0}$) of a constant size of $c$ qubits s.t. $c<n/2$ which can be created in constant depth through the application of $c$ $CZ$ gates. We claim that the size of the GHZ state can be doubled in constant depth. This can be done as follows.
@@ VI-A Log depth GHZ state generation
    $\forall q_{i}:1\leq i\leq c\implies CZ(q_{i},q_{c+i})$. The gates can be carried out in parallel, taking one time step.
@@
    $\forall q_{i}:1\leq i\leq c\implies LC(q_{i})$. Also takes one time step.
@@
    $\forall q_{i}:1\leq i\leq c\implies CZ(q_{i},q_{c+i})$. This is identical to step 1, thus it also takes one time step.
The above protocol is equivalent to applying $CZ(q_{0},q_{c+i}),\forall q_{i}:1\leq i\leq c$ but only takes three time steps (instead of $c$ time steps). Thus, the size of the GHZ state is doubled from $c$ to $2c$ in constant time. Repeated application of this protocol clearly leads to the creation of an $n$-qubit GHZ state in $O(\log(n))$ depth.
VI-B Application to subgraph complementation
Now, the above protocol can be utilized to reduce the times for steps 1 and 3 of the subgraph complementation protocol from Section III-A. We first observe that in both of those steps, a star graph is being added (modulo 2) on the induced subgraph of the qubits $\forall i:a_{i}$ with $a_{0}$ as the center.
Let there be $n$ qubits in the central node. Naively, it takes $n-1$ $CZ$ gates (and time steps) to add the star graph. We claim that by using $k$ auxiliary qubits $q_{0},q_{1}...q_{k-1}$ in the central node, the star graph can be added in $O(n/k+\log(k))$ time steps. Since steps 2 and 4 of the subgraph complementation protocol both take one time step, this reduces the asymptotic runtime for carrying out an $SC$ operation to the time required to add a star graph (in steps 1 and 3). The protocol is described as follows.
Figure 4: Parallelized subgraph complementation. The solid lines indicate the inner edges whereas the dashed lines indicate the outer edges which were added in step (c) of the parallelized $SC$ operation protocol.
Parallelized SC operation protocol for steps 1 and 3
@@ VI-B Application to subgraph complementation
    Create a $k$-qubit GHZ state at the central node on the auxiliary $k$ qubits by using the protocol from Section VI-A. This takes $O(\log(k))$ time steps.
@@
    Add $(k-1)$ edges of the form $(q_{0},a_{i})$ through the doubling protocol described in section VI-A . This can be done on any subset of size $(k-1)$ of qubits $a_{i}$.
@@
    Repeat step (b) $n/(k-1)$ times to create a star graph with $(k-1)$ inner edges (from the initial $k$-qubit GHZ state) and $n$ outer edges (between $q_{0}$ and $a_{i}$). Fig. 4 illustrates this construction.
@@
    Remove the $(k-1)$ inner edges, similar to step (a) This also takes $O(\log(k))$ time steps.
Note that step 4 of the original subgraph complementation protocol now becomes $LC(q_{0})$ instead of $LC(a_{0})$.
VI-C Analysis
Let $T(G)$ denote the time required by the parallelized protocol to distribute a graph state $\left|G\right\rangle$. Denote by $t(n)$ the time required for a single subgraph complementation on $n$ vertices, $n\in\mathbb{N}$. Steps 1 and 4 of the above protocol take $O(\log n)$ time. Steps 2 and 3 take $O(n/k)$ time. The other steps in the original subgraph complementation protocol take constant time. Thus, with $(n+k)$ qubits in the central node,
@@ VI-C Analysis
We can show that a tighter bound holds with high probability for the distribution of random graphs. The minimum rank of random graph over the field $\mathbb{F}_{2}$ is bounded according to the following theorem [18].
Theorem VI.1.
For an Erdős-Rényi random graph $\mathcal{G}_{n,p}$, and any constant $10<p<$ 1, it holds with high probability that
$\mathbb{c}_{2}(\mathcal{G}_{n,p})=\Theta\left(\frac{n}{\log n}\right).$
Thus, since $t(n)=O(\log n)$ by the prior result, we have with high probability that
$T(\mathcal{G}_{n,p})=\mathbb{c}_{2}(\mathcal{G}_{n,p})\cdot t(\overline{n})=O\left(\frac{n}{\log n}\right)\cdot O(\log n)=O(n)\,.$ (4)
for a random graph $\mathcal{G}_{n,p}$.
VII Graph State Generation in the Presence of Noise
In this section, we present numerical results comparing our graph state distribution algorithm to the algorithm proposed by Avis et al. [8], which is closest to ours in terms of problem setting and resource consumption  [^1] All the code used to carry out simulations, including programs to create the graphs in this sections, are publicly available at https://github.com/Anisen123/graph-state-simulator..
@@ VII Graph State Generation in the Presence of Noise
This is equivalent to applying $X,Y$ or $Z$ gates on $\rho$ with equal probability $p/3$ and the identity $I$ with probability $(1-p)$. Each time an operation ($LC$ or $CZ$) is carried out on some qubit $q_{i}$ with state $\rho_{i}$, we model noise by applying 5 on the qubit. Thus, the graph state $\left|G\right\rangle$ is transformed as follows:
@@
for some bitstring $b=b_{1}b_{2}...b_{n}$, where $Z$ is applied on the $i$’th qubit iff $b_{i}=1$, since the action of any Pauli operator on $q_{i}$ can be mapped to $Z$ gates acting on some set of qubits in the graph state [20]. Thus, we can now express all noise acting on $\left|G\right\rangle$ as a single bitstring of $Z$ operators acting on $\left|G\right\rangle$ since $Z$ gates commute. This makes the noise easier to characterize and store in a list data structure. We estimate the fidelity by running this simulation several times (between 5000 and 10000 times) and finding the fraction of states where $b$ is the all zero bitstring. Note that the probability that $b$ is the all zero bitstring corresponds to the fidelity of the total mixed state. This is because the set of all possible states of the form $Z^{b}\left|G\right\rangle$ corresponding to $2^{n}$ possible linearly independent $Z$-bitstrings are all orthogonal to each other.
@@
VII-A GHZ State Generation
Fig. 5 compares the fidelity obtained using our protocol with that obtained using the protocol by Avis et al. [8] to generate a 3-qubit GHZ state, as a function of probability $p$ of applying a non-identity Pauli gate at each operation. We note that we implement the local creation of a GHZ state (for [8]) by creating a star graph state first and then applying an $LC$ operation on the central node. Fig. 5 shows that our algorithm distributes small GHZ states with higher fidelity. Fig. 6 compares fidelities varying the number of qubits for a fixed $p=10^{-4}$, and we observe that there is a larger difference in fidelity to distribute graph states with a higher number of qubits.
Figure 5: Comparison of fidelity of a 3-qubit GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 6: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
VII-B Complete bipartite graph state generation
Complete bipartite graphs have applications in entanglement purification [21], silicon-photonic quantum computing [22] and are also used as graph states for all-photonic repeaters [12]. Here, we consider complete bipartite graphs with equal size partitions. Fig. 7(a) and Fig. 7(b) compare the fidelity obtained by implementing our algorithm with the algorithm of Avis et al. [8] for complete bipartite graph states with 6 qubits and 20 qubits respectively. Our algorithm creates complete bipartite graph states with lower fidelity when the number of qubits is small (less than 6 qubits), but with higher fidelity for larger states with 20 qubits. Fig. 8 compares fidelity when the number of qubits is varied between 6 and 30 (with steps of 2 qubits), for error probability $p=10^{-3}$.

### v5

@@ 4.2 Time for subgraph complementation
[…]
Given a particular $\mathcal{SC}_{sys}$, we can also optimize the ordering. Note that a complete edge reset is not always necessary; some edges can potentially be preserved, thereby reducing both the number of required gates and the total number of time steps.
@@
If $c_{2}(G)$ is constant with respect to some class of graphs, then the graph states corresponding to that class can be efficiently distributed in a linear number of time steps since $t(\overline{k})\leq n$ so $T(G,\mathcal{SC}_{sys})=O(n)$ (from 4.1). It turns out that we can characterize the class of graphs with a constant bounded value of $c_{2}(G)$ [5]. Dense structured graphs often have a bounded $\mathbb{c}_{2}(G)$, such as cliques ($\mathbb{c}_{2}(G)=1$) and complete bipartite graphs $(\mathbb{c}_{2}(G)=3)$. Specifically, the minimum distribution time $T(G,\mathcal{SC}_{sys})$ for a clique is achieved with $\mathcal{SC}_{sys}=\{V\}$. Similarly, for complete bipartite graphs with vertex partitions $P_{1},P_{2}$ the corresponding optimal subgraph complementation system is given by $\mathcal{SC}_{sys}=\{P_{1}\cup P_{2},P_{1},P_{2}\}$.
@@
[…]
Table 1: Costs for distributing various graph states
@@
[…]
Path Graph n1n-1 $O(1)$ $O(n)$
Cycle Graph n2n-2 $O(1)$ $O(n)$
@@
[…]
Complete Graph ($n$-GHZ state) 11 $O(n)$ $O(n)$
Repeater Graph [3] $\leq 3+n$ $O(1)$ $O(n)$
@@
4.3 Entanglement of the graph state
An interesting property of $\mathbb{c}_{2}(G)$ is its relation to the Schmidt-rank $r_{G}(X)$, a common entanglement measure of a pure state. For a graph state $\left|G\right\rangle$, $r_{G}(X)$ is defined over some bipartition $(X,V(G)\setminus X)$ of the vertices of $G$. For a pure state, the Schmidt-rank counts the minimal number of superpositions of factorizable states needed to generate the state [22] – a higher Schmidt-rank of a state implies it has higher entanglement. Define the average Schmidt-rank $\mathbb{E}_{r}(G)$ to be the expectation of $r_{G}$ for a uniformly chosen random subset of $V$.
Theorem 4.3.
@@ 4.3 Entanglement of the graph state
Consequently, graph states with a high degree of entanglement will have large values of the subgraph complementation number $\mathbb{c}_{2}(G)$, and therefore likely to be more expensive to distribute.
5 Detailed comparison to prior work
Our approach utilizes (n1)(n-1) EPR pairs, $O(\overline{k}n)$ bits of classical communication (CC), and a maximum of n1n-1 qubits at the central node, all of which improve on prior work [16, 12, 2, 10]. Our approach requires one time step to distribute an arbitrary graph state, if we follow early work to implement our cost model [16, 4], which make the assumption that local gates at multiple nodes and the same Clifford operations on a single central node commute, so all the operations can be performed in one time step. We instead assume a gate-based model of computation, which implies $T(G,\mathcal{SC}_{sys})$ number of time steps is required. Other work [2, 10] utilizes an approach with a central network node that takes $O(n)$ time steps if local operations are free and $O(|E|)=O(n^{2})$ time steps if not (accounting for local gates) to build a graph state corresponding to a dense graph. In the presence of memory decoherence, local operations are not free. The fidelity of the state directly depends on the number of time steps required. Our approach also utilizes $O(n^{2})$ number of gates and time steps to distribute certain dense graph states. However, many common graph states, e.g., GHZ states, utilize a number of gates and time steps that is $O(n)$ since $T(G,\mathcal{SC}_{sys})=O(n)$ for those graph states. This presents a quadratic improvement in the number of time steps, gates required, and potential fidelity of the state in the presence of noise. Table 4.2 includes the various costs of different classes of graphs.
6 Incorporating Parallelization
We first describe a log depth circuit to generate a $n$ qubit GHZ state, that is the same up to local rotations as the construction proposed by Cruz et al. [9]. This will be useful as a subroutine in reducing the time to create a star graph, a graph consisting of one central vertex connected with edges to all the other vertices, as in steps 1 and 3 of the subgraph complementation operation from 3.1.
6.1 Log depth GHZ state generation
Denote the $n$ qubits as $q_{0},q_{1},q_{2}...q_{n-1}$. Begin with a GHZ state (star graph state with center $q_{0}$) of a constant size of $c$ qubits s.t. $c<n/2$ which can be created in constant depth through the application of $c$ CZCZ gates. We now present the procedure to double the size of the GHZ state in constant depth.
GHZ state doubling algorithm
@@ 6.1 Log depth GHZ state generation
[…]
@@
    For $i=1$ to $c$ in parallel, apply $CZ(q_{i},q_{c+i})$.
@@
    For $i=1$ to $c$ in parallel, apply $LC(q_{i})$.
  4. 4.
    For $i=1$ to $c$ in parallel, apply $CZ(q_{i},q_{c+i})$.
  5. 5.
[…]
Time Analysis: Each step in the GHZ state doubling algorithm consists of $c$ operations that can be performed in parallel. Thus, each for-loop contributes only one time step, and the entire procedure takes constant depth (three time steps in total). The above algorithm is equivalent to applying $CZ(q_{0},q_{c+i}),\forall q_{i}:1\leq i\leq c$ but only requires three time steps (instead of $c$ time steps). Thus, the size of the GHZ state is doubled from $c$ to $2c$ in constant time. Repeated application of this algorithm clearly leads to the creation of an $n$-qubit GHZ state in $O(\log(n))$ depth.
6.2 Application to subgraph complementation
The above algorithm can be utilized to reduce the times for steps 1 and 3 of the subgraph complementation operation (from 3.1). We first observe that both steps induce a star graph (with the edges added modulo 2) on qubits $a_{0},\ldots,a_{k-1}$ with $a_{0}$ as the center. Let $n$ qubits reside in the central node. Naively, n1n-1 CZCZ gates (and time steps) are required to construct the star graph. We claim that by introducing $k$ auxiliary qubits $q_{0},q_{1}...q_{k-1}$, $k=1,2,\ldots,n$, in the central node, the star graph can be generated in $O(n/k+\log(k))$ time steps. Since steps 2 and 4 of the subgraph complementation operation both take one time step, this reduces the asymptotic runtime for carrying out an $SC$ operation to the time required to add a star graph (in steps 1 and 3). The parallelized SC operation is described as follows.
Figure 4: Parallelized subgraph complementation. The solid lines indicate the inner edges whereas the dashed lines indicate the outer edges which were added in step (c) of the parallelized SCSC operation.
Parallelized SC operation for steps 1 and 3
@@ 6.2 Application to subgraph complementation
    Consider a network with a central node $A$ that contains $n$ primary qubits $a_{0},...,a_{n-1}$ and $k$ auxiliary qubits $q_{0},...,q_{k-1}$. The network also contains n1n-1 nodes $C_{0},...,C_{n-2}$ each containing a qubit $c_{0},...,c_{n-2}$.
@@
    Create a $k$-qubit GHZ state at the central node on the auxiliary qubits $q_{0},...,q_{k-1}$ by repeating the GHZ state doubling algorithm $\log(k)$ times (from 6.1).
@@
    Add (k1)(k-1) edges of the form $(q_{0},a_{i})$ through the GHZ state doubling algorithm (from 6.1) . This can be done on any subset of size (k1)(k-1) of qubits $a_{i}$.
@@
    Repeat step (b) n/(k1)n/(k-1) times to create a star graph with (k1)(k-1) inner edges (between $q_{0}$ and $q_{i}:i\neq 0$, from the initial $k$-qubit GHZ state) and $n$ outer edges (between $q_{0}$ and $a_{i}$). Fig. 4 illustrates this construction.
  5. (e)
    Remove the (k1)(k-1) inner edges, similar to step (a).
Note that step 4 of the original subgraph complementation operation (from 3.1) now becomes $LC(q_{0})$ instead of $LC(a_{0})$.
6.3 Analysis
Let $T(G)$ denote the time required by the parallelized SC algorithm to distribute a graph state $\left|G\right\rangle$. Denote by $t(n)$ the time required for a single subgraph complementation on $n$ vertices, $n\in\mathbb{N}$. Steps (b) and (e) of the above operation take $O(\log n)$ time. Steps (c) and (d) take $O(n/k)$ time. The other steps in the original subgraph complementation operation (from 3.1) take constant time. Thus, with $(n+k)$ qubits in the central node,
@@ 6.3 Analysis
This derivation has the simple corollary that is if $\mathbb{c}_{2}(G)=m$ for some constant $m$, then $T(G)=O(\log(n))$. In particular, this provides an exponential improvement for generating classes of graph states corresponding to such graphs when compared to direct parallelization of applying CZ gates, which is the best known method generally, and which requires $O(n)$ circuit depth [7]. An example of such a class of graphs is a complete $m$-partite graph, which has $\mathbb{c}_{2}(G)=m$ for some constant $m$. We discuss this further in section 7.3.
7 Graph State Generation in the Presence of Noise
In this section, we present numerical results comparing our graph state distribution algorithm to the algorithms proposed by Avis et al. [2] and Cuquet, Calsamiglia [10], which are closest to ours in terms of problem setting and resource consumption [^1] All the code used to carry out simulations, including programs to create the graphs in this section, are publicly available at https://github.com/Anisen123/graph-state-simulator..
@@ 7 Graph State Generation in the Presence of Noise
This is equivalent to applying $X,Y$ or $Z$ gates on $\rho$ with equal probability $p/3$ and the identity $I$ with probability (1p)(1-p). Each time an operation ($LC$ or $CZ$) is carried out on some qubit(s) $q_{i}$ with state $\rho_{i}$, we model noise by applying 4 to each qubit. Thus, the graph state $\left|G\right\rangle$ is transformed as follows:
@@
[…]
@@
7.1 GHZ State Generation
Noise Model. We consider noise incurred from carrying out $CZ,LC$ gate operations, measuring qubits in the $Y$ or $Z$ basis, and local Clifford correction unitaries. We assume Bell pairs that are initially noiseless. While noisy Bell pairs will impact the final fidelity, we do not expect the relative performance of the algorithms to change, since each algorithm uses roughly the same initial number of Bell pairs (in fact, our algorithm uses one less Bell pair). In Appendix A, we provide simulation results that consider memory decoherence as well in Figs. 13,14. First we compare the fidelity obtained using our algorithm (from 4.1) with that obtained using the Factory Node + Teleportation (FNT) Algorithm (from 1.1) (or the Bipartite A protocol) and the Bipartite B protocol [10] to generate a 3-qubit GHZ state. Fig. 5 (left) compares the fidelity as a function of probability $p$ of applying a non-identity Pauli gate at each operation. Note that we implement the local creation of a GHZ state (for Bipartite A and B) by creating a star graph state first and then applying an $LC$ operation on the central node.
Figure 5: Comparison of fidelity of a 3-qubit GHZ state distributed according to the SC algorithm (4.1), the FNT algorithm (Bipartite A protocol) (1.1), and the Bipartite B protocol, considering (left) and neglecting (right) noise from local correction unitaries, having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
We observe that our algorithm generates higher fidelities than the others. We also note that the Bipartite A protocol generates GHZ states with higher fidelity compared to the Bipartite B protocol, despite the latter using fewer gate operations [10]. Note that this differs from prior results reported by Cuquet and Calsamiglia [10]. However, we recover their results in Fig. 5 (right), when we neglect the noise from local correction unitaries, as was done in their paper but which we do not. We support these results by showing that the same trend holds when generating graph states with a varying number of qubits, Fig. 11 in Appendix A. Consequently, to simplify analysis, we compare our algorithm only with the FNT Algorithm (1.1) (or Bipartite A) for the rest of the paper since it experimentally performs better following our noise model.
Fig. 6 shows fidelity vs. number of qubits for a fixed $p=10^{-4}$ for both the FNT and Subgraph Complementation (SC) Algorithms. We observe little or no difference between these two algorithms when the number of qubits is less than or equal to $x$, and that the SC algorithm produces higher fidelity states otherwise and that the difference increases with number of qubits.
Figure 6: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the SC algorithm (4.1) and the FNT algorithm [2] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
7.2 Complete bipartite graph state generation
Complete bipartite graphs have applications in entanglement purification [23], silicon-photonic quantum computing [20] and are also used as graph states for all-photonic repeaters [3]. Here, we consider complete bipartite graphs with equal size partitions. Fig. 6(a) and Fig. 6(b) compare the fidelity obtained by our algorithm with that of the FNT algorithm (1.1) [2, 10], for complete bipartite graph states with 6 qubits and 20 qubits respectively. Our algorithm creates complete bipartite graph states with lower fidelity when the number of qubits is small (less than 12 qubits), but with higher fidelity for larger states with 20 qubits. Fig. 8 compares fidelity when the number of qubits is varied between 6 and 30 (with steps of 2 qubits), for error probability $p=10^{-3}$.